Solve the system algebraically.
2x + y - 10 = 0
x - y - 5 = 0
What is the value of y?

Answer by mananth(16946)

2 x + 1 y = 10 .............1
1 x -1 y = 5 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 1
2 x + 1 y = 10
1 x + -1 y = 5
Add the two equations
3 x = 15
/ 3
x = 5
plug value of x in (1)
2 x + 1 y = 10
10 + 1 y = 10
1 y = 10 -10
y = 0